OpenStudy (anonymous):

HELP PLEASE (x+2)^2/9 + (y-3)^2/16 =1 Can Someone Help me Figure out what kind of conic section it is?

OpenStudy (anonymous):

\[\frac{ (x+2)^{2} }{ 9 } + \frac{ (y-3)^{2} }{ 16} = 1\]

OpenStudy (amoodarya):

vertical ellipsoid

OpenStudy (anonymous):

Thank you :)